Circular No. 7739

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2001 RX_{14} (LINEAR) The linkage by the Minor Planet Center of subsequent observations of

an apparent main-belt minor planet observed by LINEAR on Sept. 10 and 11 showed that this object, designated 2001 RX_{14} on MPS 34978, had a nearly parabolic orbit. After placement on the NEO Confirmation Page, many further observations were received, with M. Tichý (Kleť, 0.57-m f/5.2 reflector + CCD) on Oct. 18.7 UT indicating that the object appeared to be diffuse with a coma of diameter 13′′ and red mag 16.7. R-band images taken by C. Hergenrother and J. Barnes (Lunar and Planetary Laboratory) with the 1.54-m Catalina reflector on Oct. 24.32 show a highly condensed coma of diameter 6′′ and no hint of a tail.

2001 UT  α_{2000}  δ_{2000}  m_{2}
Sept.10.32152  2^{h}02^{m}43.^{s}50  +22^{o}36′57′′.9  18.8

The available astrometry, including prediscovery observations made by the NEAT and LONEOS programs on Aug. 28, along with the following parabolic orbital elements, appear on MPEC 2001-U46.

T = 2003 Jan. 19.036 TT  ω = 121.^{o}496
Ω = 14.171 }2000.0
q = 2.05903 AU  i = 30.560
2001TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Oct. 18 1^{h}42.^{m}56 +23^{o}15.^{′}1 4.273 5.246 166.^{o}0 2.^{o}6 16.9
28 1 35.55 +23 08.5 4.189 5.163 167.1 2.5 16.7
Nov. 7 1 28.59 +22 56.3 4.137 5.080 160.0 3.8 16.6
17 1 22.11 +22 40.5 4.115 4.996 150.1 5.7 16.6
27 1 16.53 +22 23.5 4.120 4.913 139.5 7.5 16.5
Dec. 7 1 12.18 +22 07.9 4.149 4.829 128.9 9.1 16.4

COMET 19P/BORRELLY Visual m_{1} estimates: Sept. 1.11 UT, 10.0 (R. J. Bouma, Groningen,

The Netherlands, 0.25-m reflector); 28.14, 9.9 (W. Hasubick, Buchloe, Germany, 25×100 binoculars); Oct. 11.08, 10.1 (B. H. Granslo, Fjellhamar, Norway, 0.20-m reflector); 18.68, 10.3 (Y. Nagai, Yamanashi, Japan, 0.32m reflector); 24.96, 10.7 (V. S. Nevski, Vitebsk, Belarus, 0.30-m reflector).

2001 October 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green